# AROMA: Preserving Spatial Structure for Latent PDE Modeling with Local Neural Fields

### v2

@@ Abstract
We present AROMA  [^1] The source code for this paper is available at https://github.com/LouisSerrano/aroma (Attentive Reduced Order Model with Attention), a framework designed to enhance the modeling of partial differential equations (PDEs) using local neural fields. Our flexible encoder-decoder architecture can obtain smooth latent representations of spatial physical fields from a variety of data types, including irregular-grid inputs and point clouds. This versatility eliminates the need for patching and allows efficient processing of diverse geometries. The sequential nature of our latent representation can be interpreted spatially and permits the use of a conditional transformer for modeling the temporal dynamics of PDEs. By employing a diffusion-based formulation, we achieve greater stability and enable longer rollouts compared to conventional MSE training. AROMA’s superior performance in simulating 1D and 2D equations underscores the efficacy of our approach in capturing complex dynamical behaviors.
@@ 4 Experiments
In this section, we systematically evaluate the performance of our proposed model across various experimental settings, focusing on its ability to handle dynamics on both regular and irregular grids. First, we investigate the dynamics on regular grids, where we benchmark our model against state-of-the-art neural operators, including Fourier Neural Operators (FNO), ResNet, Neural Fields, and Transformers. This comparison highlights the efficacy of our approach in capturing complex spatio-temporal patterns on structured domains. Second, we extend our analysis to dynamics on irregular grids and shared geometries, emphasizing the model’s extrapolation capabilities in data-constrained regimes. Here, we compare our results with Neural Fields and Transformers, demonstrating the robustness of our model in handling less structured and more complex spatial configurations. Lastly, we assess the model’s capacity to process diverse geometries and underlying spatial representations by comparing its performance on irregular grids and different geometries. This evaluation highlights the flexibility and generalization ability of our model in encoding and learning from varied spatial domains, showcasing its potential in accurately representing and predicting dynamics across a wide range of geometric settings.
@@ 6 Conclusion and Limitations
AROMA offers a novel principled and flexible neural operator approach for modeling the spatio-temporal evolution of physical processes. It is able to deal with general geometries and to forecast at any position of the spatial domain. It incorporates in a principled framework attention mechanisms, a latent diffusion transformer for spatio-temporal dynamics and neural fields for decoding. Thanks to a very compact spatial encoding, its complexity is lower than most SOTA models. Experiments with small-size datasets demonstrate its effectiveness. Its reduced complexity holds potential for effective scaling to larger datasets. As for the limitations, the performance of AROMA are still to be demonstrated on larger and real world examples. Also, even though it has some potential for uncertainty modeling, this aspect has still to be further explored and analyzed.
@@ C.2 Encoding interpretation
C.3 Long rollouts
We show long rollout predictions using AROMA on Burgers dataset in Figure 10, and on Navier-Stokes $1\times 10^{-3}$ dataset in Figure 12. AROMA returns predictions that remain stable and accurate, even outside the training time horizon.
@@ C.3 Long rollouts
Figure 11: Uncertainty of AROMA over rollout steps
@@
Figure 12: Test example rollout trajectories with AROMA on Navier-Stokes. Top: predicted trajectory on In-t. Bottom: trajectory on Out-t. First row in each subfigure shows the prediction, the second row shows the ground truth.
@@ C.5 Spatial tokens perturbation analysis
To validate the spatial interpretation of our latent tokens, we establish a baseline code ${\bm{Z}}^{0}$, and introduce perturbations by sequentially replacing the $j$-th token, ${\bm{z}}^{0}_{j}$, with subsequent tokens along the trajectory, denoted as ${\bm{z}}^{1}_{j},{\bm{z}}^{2}_{j},\ldots,{\bm{z}}^{t}_{j}$. Thus, the perturbed tokens mirror ${\bm{Z}}^{0}$ in all aspects except for the $j$-th token, which evolves according to the true token dynamics. We show reconstruction visualizations of the perturbed tokens in figs. 14, 15, 16, 17, 18, 19, 20 and 21. On the right side, we show the groundtruth of the trajectory. On the left side, is the change in AROMA’s prediction in response to the token perturbation. These figures show that the perturbation of a token only impacts the reconstructed field locally, which validates the spatial structure of our tokens. Additionally, we can notice some interesting effects of the token perturbations near the boundaries in figs. 16 and 21: our encoder-decoder has discovered from data and without explicit supervision that the solutions had periodic boundary conditions by leveraging the encoded geometry and the function values. This validates the architecture of our cross-attention module between the function values, the spatial coordinates and the geometry-aware tokens.
@@
C.6 Ablation study
We show the impact of the number of latent tokens on the Navier-Stokes1e-4 dataset in Table 6. We train our auto-encoder with 10000 trajectories. We can see that the performance increases with the number of tokens.
@@ C.6 Ablation study
Table 6: Influence of the number of latent tokens on the test reconstruction capabilities on NavierStokes1e-4. Performance in Relative $L_{2}$ Error.
C.7 Latent space dynamics
For Navier-Stokes, we show how the mean (Figure 22) and standard deviation tokens (Figure 23) evolve over time for a given test trajectory. We show the predicted trajectory of the latent tokens $Z$ in the latent space in Figure 24.
In practice, the tokens where the logvar is 00 on Figure 23 do not impact the prediction (Rolinek et al., 2019).

### v3

@@ Abstract
We present AROMA (Attentive Reduced Order Model with Attention), a framework designed to enhance the modeling of partial differential equations (PDEs) using local neural fields. Our flexible encoder-decoder architecture can obtain smooth latent representations of spatial physical fields from a variety of data types, including irregular-grid inputs and point clouds. This versatility eliminates the need for patching and allows efficient processing of diverse geometries. The sequential nature of our latent representation can be interpreted spatially and permits the use of a conditional transformer for modeling the temporal dynamics of PDEs. By employing a diffusion-based formulation, we achieve greater stability and enable longer rollouts compared to conventional MSE training. AROMA’s superior performance in simulating 1D and 2D equations underscores the efficacy of our approach in capturing complex dynamical behaviors. Github page: https://github.com/LouisSerrano/aroma
@@ 4 Experiments
In this section, we systematically evaluate the performance of our proposed model across various experimental settings, focusing on its ability to handle dynamics on both regular and irregular grids. First, we investigate the dynamics on regular grids, where we benchmark our model against state-of-the-art neural operators, including Fourier Neural Operators (FNO), ResNet, Neural Fields, and Transformers. This comparison highlights the efficacy of our approach in capturing complex spatio-temporal patterns on structured domains. Second, we extend our analysis to dynamics on irregular grids and shared geometries, emphasizing the model’s extrapolation capabilities in data-constrained regimes. Here, we compare our results with Neural Fields and Transformers, demonstrating the robustness of our model in handling less structured and more complex spatial configurations. Lastly, we assess the model’s capacity to process diverse geometries and underlying spatial representations by comparing its performance on irregular grids and different geometries. This evaluation highlights the flexibility and generalization ability of our model in encoding and learning from varied spatial domains, showcasing its potential in accurately representing and predicting dynamics across a wide range of geometric settings. We include additional results from ablation studies in Section C.6.
@@ 6 Conclusion and Limitations
AROMA offers a novel and flexible neural operator approach for modeling the spatio-temporal evolution of physical processes. It is able to deal with general geometries and to forecast at any position of the spatial domain. It incorporates in an encode-process-decode framework attention mechanisms, a latent diffusion transformer for spatio-temporal dynamics and neural fields for decoding. Thanks to a very compact spatial encoding, its complexity is lower than most SOTA models. Experiments with small-size datasets demonstrate its effectiveness. Its reduced complexity holds potential for effective scaling to larger datasets. As for the limitations, the performance of AROMA are still to be demonstrated on larger and real world examples. Moreover, like all dynamical models that operate over a latent space, the reconstruction capabilities of the decoder is a bottleneck for the rollout accuracy. Since the encoder and decoder are learning spatial relationships from scratch, conferring the framework a high flexibility, the training efficiency does not match that of CNN-based auto-encoders on regular grids. We therefore believe there could be further improvements to be made to achieve a similar performance while keeping the same level of flexibility. Finally, even though our model has some potential for uncertainty modeling, this aspect has still to be further explored and analyzed.
Acknowledgements
We acknowledge the financial support provided by DL4CLIM (ANR-19-CHIA-0018-01), DEEPNUM (ANR-21-CE23-0017-02), PHLUSIM (ANR-23-CE23-0025-02), and PEPR Sharp (ANR-23-PEIA-0008, ANR, FRANCE 2030).
@@ C.2 Encoding interpretation
C.3 Example rollouts
We show examples of rollout predictions using AROMA on Burgers dataset in Figure 10, on Navier-Stokes $1\times 10^{-3}$ dataset in Figure 12 and on CylinderFlow in Figure 13. AROMA returns predictions that remain stable and accurate, even outside the training time horizon.
@@ C.3 Example rollouts
Figure 11: Uncertainty of AROMA over rollout steps. The blue line is the mean prediction while the blue shade represents the $\text{mean}\pm 3\times\text{standard deviation}$.
@@
Figure 12: Test example rollout trajectories with AROMA on Navier-Stokes $1\times 10^{-3}$. Top: predicted trajectory on In-t. Bottom: trajectory on Out-t. First row in each subfigure shows the prediction, the second row shows the ground truth.
Figure 13: Visualization of AROMA’s predictions on Cylinder for (Out-t). The left panel shows the prediction, the middle panel displays the ground truth, and the right panel is the point-wise error.
@@ C.5 Spatial tokens perturbation analysis
To validate the spatial interpretation of our latent tokens, we establish a baseline code ${\bm{Z}}^{0}$, and introduce perturbations by sequentially replacing the $j$-th token, ${\bm{z}}^{0}_{j}$, with subsequent tokens along the trajectory, denoted as ${\bm{z}}^{1}_{j},{\bm{z}}^{2}_{j},\ldots,{\bm{z}}^{t}_{j}$. Thus, the perturbed tokens mirror ${\bm{Z}}^{0}$ in all aspects except for the $j$-th token, which evolves according to the true token dynamics. We show reconstruction visualizations of the perturbed tokens in figs. 15, 16, 17, 18, 19, 20, 21 and 22. On the right side, we show the groundtruth of the trajectory. On the left side, is the change in AROMA’s prediction in response to the token perturbation. These figures show that the perturbation of a token only impacts the reconstructed field locally, which validates the spatial structure of our tokens. Additionally, we can notice some interesting effects of the token perturbations near the boundaries in figs. 17 and 22: our encoder-decoder has discovered from data and without explicit supervision that the solutions had periodic boundary conditions by leveraging the encoded geometry and the function values. This validates the architecture of our cross-attention module between the function values, the spatial coordinates and the geometry-aware tokens.
@@
C.6 Ablation studies
Number of tokens $M$
We show the impact of the number of latent tokens on the Navier-Stokes $1\times 10^{-4}$ dataset in Table 6. We train our auto-encoder with 10000 trajectories. We can see that the performance increases with the number of tokens.
@@ C.6 Ablation studies
Table 6: Influence of the number of latent tokens on the test reconstruction capabilities on Navier-Stokes $1\times 10^{-4}$. Performance in Relative $L_{2}$ Error.
Auto-encoding vs VAE
Our framework can also be used without the KL regularization, and could potentially be employed with other forms of regularization, such as L2 regularizaton or vector-quantization (Oord et al., 2017). We investigated in Table 7 the impact the KL regularization had on the overall rollout performance, and selected an autoencoder with L2 regularization (weight decay) as baseline. Our conclusion is that using an autoencoder with L2 regularization is a viable alternative to the VAE in some cases for achieving a smooth latent space. The autoencoder demonstrated superior performance on two datasets (Burgers and Navier-Stokes $1\times 10^{-4}$, explained by its lower reconstruction errors, which translate into better rollout performance. However, for the more challenging Navier-Stokes $1\times 10^{-5}$ case, the autoencoder’s latent space exhibited high variance, which may explain the observed performance difference with the VAE.
No-diffusion vs diffusion
As an ablation, we also measured the influence of the diffusion formulation on the rollout accuracy by comparing to the same transformer architecture trained directly with an MSE on the mean tokens. The deterministic version of AROMA shows consistently robust performance and even surpasses the diffusion version on the Navier-Stokes $1\times 10^{-4}$ case (Table 7). This demonstrates that the latent tokens obtained with AROMA contain meaningful information for dynamics modeling. On the other hand, the deterministic version yields less accurate long rollouts on Burgers or KS in Figure 23 and Figure 3. Note that using diffusion allows us to model the trajectory distribution, which opens the way to infer statistics on this distribution. This is key, for example, when modeling uncertainty, which is a critical problem for these models.
Latent MLP vs Latent Transformer
Modeling interactions at the local and global levels is key to learn the dynamics faithfully. Experiments using MLPs (Table 7) as time steppers which do not consider interactions between tokens lead to significantly lower performance compared to transformers.
Table 7: Ablation Study. Metrics in Relative $L_{2}$ on the test set.
Model Burgers Navier-Stokes Navier-Stokes
$1\times 10^{-4}$ $1\times 10^{-5}$
AROMA + auto-encoding $\mathbf{3.43\times 10^{-2}}$ $\mathbf{5.02\times 10^{-2}}$ $2.10\times 10^{-1}$
AROMA w/o diffusion ${4.31\times 10^{-2}}$ ${7.50\times 10^{-2}}$ ${1.28\times 10^{-1}}$
AROMA + mlp ${1.11\times 10^{-1}}$ ${1.00\times 10^{0}}$ ${8.25\times 10^{-1}}$
AROMA ${3.65\times 10^{-2}}$ ${1.05\times 10^{-1}}$ $\mathbf{1.24\times 10^{-1}}$
C.7 Kuramoto-Sivashinsky : a failure case
We conducted additional experiments on a chaotic 1D PDE, the Kuramoto-Sivashinsky (KS) equation. We found that AROMA currently struggles with dynamics that exhibit chaotic phenomena and non-decaying spectra, as shown in Figure 23. The primary limitation appears to be the reconstruction capabilities of the encoder-decoder. For the KS equation, we found that obtaining reconstructions with an MSE in the range of 1e-10 to 1e-12 was necessary for accurate spectrum reconstruction. Like all models leveraging a reduced latent representation space, AROMA inherently loses some of the fine-grained details necessary for accurately capturing chaotic behavior. The diffusion framework slightly improves the high correlation time compared to the deterministic version, however the main bottleneck comes from the decoder (Table 8). In conclusion, while AROMA performs very well on simpler dynamics, dealing with chaotic phenomena requires more involved modeling that explicitly targets the chaotic component. Note that using dedicated modules for this purpose is a current practice in fluid dynamics - e.g. LES (Large eddy simulation).
Table 8: Test results on the KS equation. The evaluated metrics include: 1-step prediction MSE, MSE over the entire rollout (160 timestamps), and the duration for which the correlation between the generated samples and the ground truth remains above 0.8.
Baseline 1 step. MSE Rollout MSE Corr. $\geq$ 0.8
AROMA w/o diffusion $\mathbf{{1.25\times 10^{-5}}}$ ${2.20\times 10^{0}}$ 32.8s
AROMA ${1.81\times 10^{-5}}$ ${\mathbf{2.07\times 10^{0}}}$ 36.0s
(a) Spectrum of AROMA’s 1-step predictions vs ground truth.
(b) Comparison of the MSE loss ($\downarrow$) with and without diffusion.
Figure 23: Qualitative results on KS equation.
C.8 Latent space dynamics
For Navier-Stokes, we show how the mean (Figure 24) and standard deviation tokens (Figure 25) evolve over time for a given test trajectory. We show the predicted trajectory of the latent tokens $Z$ in the latent space in Figure 26. In practice, the tokens where the logvar is 00 (i.e. a high variance) on Figure 25 do not impact the prediction (Rolinek et al., 2019). We can therefore see, that ouf of the 16 tokens used, the most influential ones are Token 6, 7, 8, 9, 15, 16, as they clearly exhibit non-noisy patterns.